Prove -(Cos30-sin20)/cos40+cos20) = (4/root3)cos80cos40

Grade:11

1 Answers

Rahul Debnath
5 years ago
cos30-sin20/cos40+cos20=cos30-cos70/cos40+cos20
=2sin50sin20/2cos30cos10=2sin50sin20/root3×cos10
=2sin50×sin(180-20)/root3×cos10=2cos40sin160/rt3×cos10
=2cos40×2cos80sin80/rt3×sin80
=4×cos40×cos80/rt3.(proved)
....I used some formulae like sin2$=2sin$cos$;cos$=sin90-$;sin$=sin180-$....;etc ....if helpful PLZ like it SIRRRRR.